5. An 18mm diameter rivet is used to fasten two 22mm thick plate. If the

shearing stress is 60 MPA. What tensile force applied each plate to shear the

bolt?


( Ans F=15.27KN )


1
Expert's answer
2022-04-01T12:53:02-0400

"P=\tau{A_{rivet}}\\
\tau=60MPA\\
A_{rivet}=\frac{\pi18^2}{4}=81\pi{mm^2}\\
P=4860\pi{N}=15.27kN\\
\text{Tensile force applied to shear the bolt is 15.27kN}"